User Tools

Site Tools


operations:log_recent_events
Date Time Experiment / Event S-band VLBI SEFD X-band VLBI SEFD Comment
Jan 17 23.00 Pointing Pointing updated with +0.155 deg El offset
Jan 18 17.00 R1466 4844 5287
Jan 20 18.30 R4466 10679 18892 No data after 01.55 UT until end due to problems with Az drive tripping out
Jan 24 17.00 R4467 5734 5231 Missed scans due to Az drives 20.12 - 22.13 and 14.00 - 14.18 UT
Jan 27 03.00 Pointing Check shows a 0.02 deg offset. Probably not significant and no model update made
Jan 31 17.00 R1468 6173 4886 Drive problems. Scans from 16.09-16.27 missed
Feb 01 17.30 T2074 DBBC freqs incorrect for VC13, VC14. Rapid drift in formatter-GPS offset. Some scans missed due to drive problems.
Feb 02 18.00 OHIG70 DBBC freqs incorrect for VC13, VC14. All scans after 15.19 lost, drive problems
Feb 03 18.30 R4468 6540 6126 Many missed scans due to drive problems
Feb 07 23.40 Pointing Offsets observed: xEL = -0.003, El = 0.02. No update to model
Feb 08 17.30 OHIG71 Frequencies OK. Multiple drive failures
Feb 17 03.10 Pointing Pointing updated with +0.05 deg El offset following tightening of encoder mounting brackets
Feb 17 PCFS Problem with Az drives tripping was fixed (software bug)
Feb 17 18.30 R4470 5458 8699
Feb 21 04.13 Pointing Offsets observed: xEL = -0.01, El = -0.015. No update to model
Feb 21 17.00 R1471
Feb 24 01.15 Pointing Offsets observed on 1921-293: xEL = -0.005, El = -0.005. No update to model. Single-dish X-band SEFD looks good
Feb 24 18.30 R4471 4425 4601
Feb 28 Pointing and calibration checks Pointing looked good, Y-factors show S-band Tsys ~110K, X-band 90K
Feb 28 17.00 R1472 Note clock jump at 02.53 UT
Mar 03 00.00 Tsys Y-factor measurements repeated: S-band Tsys ~110K, X-band 85K
Mar 03 18.30 R4472
Mar 06 23.00 Antenna Checked feed for water: appears dry. Tightened some SMA connectors (a few loose ones), sealed around S-band loop couplers, checked bolts on El encoder mounting
Mar 07 01.00 Tsys Y-factors measured S-band Tsys 95-100K, X-band 83 K
Mar 07 10.00 Pointing Pointing looks good, both axes
Mar 07 17.00 R1473
Mar 15 04.00 Pointing Pointing offsets xEL = 0.008, El = -0.02. No model change made but need to watch El
Mar 15 17.00 R1474 Data shipped
Mar 17 18.30 R4474 Data shipped
Mar 21 17.00 R1475 Data shipped
Mar 22 17.30 T2075 Data shipped
Mar 24 R4475 Data shipped by eTransfer, log sent
Mar 28 Pointing Pointing looks good
Mar 28 RD1102 Data shipped, log sent
Mar 29 CRDS50 Data lost?
Mar 30 R4476 Data eTransfered, log sent
Apr 4 R1477 Module shipped, log sent
Apr 5 RDV86 Module not sent, log sent
Apr 6 CRDS51 Clock drift?
Apr 7 R4477
Apr 11 Pointing Pointing looks good
Apr 11 R1478
Apr 14 R4478
Apr 18 R1479
/home/www/auscope/opswiki/data/pages/operations/log_recent_events.txt · Last modified: 2011/10/26 06:37 (external edit)